OpenStudy (anonymous):

Find the limits of the ff. functions. If the limit does not exist, specify if it is positive or negative infinity. If it's neither, just write DNE. PLEASE HELP

OpenStudy (anonymous):

\[1. \lim_{x \rightarrow +7} \frac{ x ^{2}+4x-3 }{ 4x-28 }\]

OpenStudy (anonymous):

oops. \[1. \lim_{x \rightarrow 7^{+}} \frac{ x ^{2}+4x-3 }{ 4x-28}\]

hartnn (hartnn):

you know whats the meaning of 7+ ??

OpenStudy (anonymous):

yup. from the right of 7?

hartnn (hartnn):

so, 4x-28 will be positive or negative ??

OpenStudy (anonymous):

uhm. that where I got confused. when i substistute 7, it'll be undefined then. but with 8, it'll be positive

hartnn (hartnn):

ok, 7+ is a number very near to 7 but GREATER than 7 so, 4x will be very near to 28 but GREATER than 28 and 4x-28 will be ..... ?

OpenStudy (anonymous):

positive? :)

hartnn (hartnn):

correct! :)

hartnn (hartnn):

and denominator is obviously positive, right ?

OpenStudy (anonymous):

so the answer here is \[+\infty\] ?

hartnn (hartnn):

so, when x-> 7+ you limit will be (a positive no.)/(number very near to 0, but positive) = + infinity yes, correct :)

hartnn (hartnn):

ask if any doubts :)

OpenStudy (anonymous):

aww, thankyou so much but i have like 6 more problems here left. can u pls help me along the way? it our take home long test :)

hartnn (hartnn):

i will try my best to help :) as long as i am online...

OpenStudy (anonymous):

thanks so much.

OpenStudy (anonymous):

here it goes \[2. \lim_{x \rightarrow -3} \frac{ 3x ^{2}+4x-15 }{ x ^{3}+27 }\]

hartnn (hartnn):

just -3 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

do i need to find first the LHL and RHL and see if they're equal? :)

hartnn (hartnn):

yesss! can u find ?

OpenStudy (anonymous):

ill try. ill get back to u in a sec. :)

hartnn (hartnn):

oh, wait no! you don't need to :P

hartnn (hartnn):

because the NUMERATOR is factorable!!

hartnn (hartnn):

can u factor the numerator and denominator ?

OpenStudy (anonymous):

aww. yeah sure. so u dont need to what? to substitute and stuff?

hartnn (hartnn):

you don't need to find LHL and RHL try to factor

OpenStudy (anonymous):

\[\frac{ (3x-5) (x+3) }{ (x+3) (x ^{2}-3x+9) }\]

OpenStudy (anonymous):

is that correct? :)

hartnn (hartnn):

absolutely! :) so what gets cancelled out ?

OpenStudy (anonymous):

x+3 :)

OpenStudy (anonymous):

can i just ask? why is LHL & RHL not necessary? i thought we should need to satisfy its left and right limits? cos it not specified in the given, i mean its just -3 :)

hartnn (hartnn):

when we can solve by directly substituting the value, LHL and RHL is not necessary :)

OpenStudy (anonymous):

ohh. i see

hartnn (hartnn):

so whatever remains, put x=-3 there ,

OpenStudy (anonymous):

oh ok. ;)

OpenStudy (anonymous):

\[- \frac{14 }{ 27 }\]?

hartnn (hartnn):

good :) its correct.

OpenStudy (anonymous):

oh, okaay. thankyou. so let me get this staright? when will i use the positive over positive thing from just like in #1? :)

hartnn (hartnn):

when you need LHL and RHL or when you are not sure about the sgns, like if x->7 and denominator has x-7 or when x->0 and the denominator has x we are not sure here, because x->y just means x is very near to y smaller or greater, don't know

OpenStudy (anonymous):

oh okay. thanks. i think here is a hard one \[\lim_{x \rightarrow -5} \frac{ x ^{2}+6x+5 }{ \sqrt{-x-1}-2 }\]

hartnn (hartnn):

do u know how to rationalize the denominator ?

OpenStudy (anonymous):

i think so. yeah. ill try. so first, i need to rationalize the denominator for ..? :)

hartnn (hartnn):

you want the reason, why we are doing this ?

OpenStudy (anonymous):

yeah :)

hartnn (hartnn):

ok, here i go. when we have f(x) = (x-3)(x+2) assume, then if we put x=3, we get a 0, right ?? which means when we get a 0, (x-3) must be a factor of f(x) similarly here, when i put x=-5 in the denominator, i get a 0!! so, x-5 MUST be a factor of denominator, i must find it by rationalizing...

OpenStudy (anonymous):

ohhhh. thanks agaiin. haha

hartnn (hartnn):

similarly for numerator, only difference is its already rationalized, just needed to be factored

OpenStudy (anonymous):

and i came up with this btw \[- (\sqrt{-x-1}+2) (x+1)\]

hartnn (hartnn):

wouldn't the denominator be -x-1 -4 ??

OpenStudy (anonymous):

yeah. but its -x-5? nd i put the negative sign outside so i could cancel out x+5?

OpenStudy (anonymous):

*and

hartnn (hartnn):

oh, you already did that! good girl :)

hartnn (hartnn):

and yes, thats correct, now you can simply plug in x=-5 :)

OpenStudy (anonymous):

Haha. :) so.. is the answer -16? :)

hartnn (hartnn):

or +16 ? O.o

OpenStudy (anonymous):

oh! haha. yeah. forgot the (-) outside. haha. my bad

OpenStudy (anonymous):

thanks for reminding. haha. overlooked it :D

OpenStudy (anonymous):

uhm, excuse me for a sec. is my realization right? regarding the LHL & RHL topic. uhm, "we can only use RHL and LHL if at first, its undefined (by using the direct substitution property at first)?

hartnn (hartnn):

see usually i first try to find the common factor, and cancel it, like in last proble, it works most of the time and there we need not use LHL and RHL unless specified.

OpenStudy (anonymous):

oh. okaay. i get it. :)

hartnn (hartnn):

but special cases when there is no common factor, and the denominator =0, then we need it, like when x->3 and denominator =x-3 and we don';t know whether x-3 will be positive or not

OpenStudy (anonymous):

ohh :)

hartnn (hartnn):

next question :)

OpenStudy (anonymous):

this kinda makes me confused. haha :) \[4. \lim_{x \rightarrow \infty} \frac{ x ^{2}-4x ^{3}+4 }{ 2+3x ^{3}-2x }\]

OpenStudy (anonymous):

i kinda get it but idk if my answer is right. :D

hartnn (hartnn):

ok, here is what i do, there are other methods too..... i substitute x = 1/y so that when x->infinity , y-> 0 so, change every x to 1/y the benefit is that, after you simplify, you will just have to put y= 0 :)

OpenStudy (anonymous):

aww, our teacher taught us a diff way tho. :)

OpenStudy (anonymous):

he said to divide the numerator and denominator by the highest degree of the denominator. is that correct too? :)

hartnn (hartnn):

did u get the answer as -4/3 ?? if yes, then correct.... yes, that method is short! but not so simple to get

hartnn (hartnn):

and ofcourse absolutely correct

OpenStudy (anonymous):

yah. -4/3 :DDD

hartnn (hartnn):

here the reason, when x->inf, 1/x -> 0 hence all of these, 1/x^2. 1/x^3, ... and so on will ->0 hence you can just put them as 0

OpenStudy (anonymous):

yup. thats exactly what he taught us :)

hartnn (hartnn):

cool :) next Q :)

OpenStudy (anonymous):

\[5. \lim_{x \rightarrow -\infty} \frac{ \sqrt{x}+x ^{2} }{ x ^{3}-5x ^{2} }\]

OpenStudy (anonymous):

idk what going to happen to \[\frac{ \sqrt{x} }{ x ^{3} }\] XD

hartnn (hartnn):

the power of denominator is simply greater than numerator so, as x->-infinity, the denominator, grows more quickly than numerator. so, final limit just = 0 (anything/ very very large = very near to 0) you will be able to appreciate this with my method of x=1/y too

OpenStudy (anonymous):

ohh. yeah. so ill just substitute 1/y to every x then? :)

hartnn (hartnn):

and for your professor's method \(\Huge \dfrac{x^m}{x^n}=x^{m-n} \\ \large \dfrac{x^{(1/2)}}{x^3 } = x^{1/2-3} \) which will also tend to 0 when x->-infinity

OpenStudy (anonymous):

ohh.

OpenStudy (anonymous):

sooo im going to use ur method then? :D

hartnn (hartnn):

whichever you are comfortable with...

hartnn (hartnn):

both will give u final answer =0

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

so whats the value for \[\frac{ \sqrt{x} }{ x^{3} }\]

hartnn (hartnn):

0

OpenStudy (anonymous):

lol. because im writing down my solutions :) thanks

OpenStudy (anonymous):

but i thought its only for \[\frac{ 1 }{ x ^{r} }\]? i mean the zero rule thing? :)

hartnn (hartnn):

yes, its is

hartnn (hartnn):

but x^1/2 / x^3 can be converted to x^r \(\Huge \dfrac{x^m}{x^n}=x^{m-n} \\ \large \dfrac{x^{(1/2)}}{x^3 } = x^{1/2-3}\) where r>1

hartnn (hartnn):

whats 1/2-3 ?

OpenStudy (anonymous):

oh. i see. so the numerator is open for any number and as long as r>1, its zero, right? :D

OpenStudy (anonymous):

-2.5? :)

hartnn (hartnn):

think like this, if exponent of numerator is LESS than exponent of denominator, then when x->infinity, it = 0

hartnn (hartnn):

x^-2.5 = 1/x^2.5 exponent >1 so 0

OpenStudy (anonymous):

oh.. i get it :D

OpenStudy (anonymous):

this is much easier to understand coming from you :D